7. Cells are kept in suspension in RPMI.a media for growing cells.You have 30 mL of a suspension at a concentration of Z2x10e cellslmL You need to dilute a portion these cells such that 100 HL the new suspension of of a 96-wel microtitre plate. will deliver 10 plated in the wells answer wells? should account How dilute cells so you could prepare Your for 5-25% extra volume as a pipetting allowance (answers may vary lo mL cells. 0.5 x lo S cells ne ed 2.2 x 10 3.1 mL RPMI (4.4 x 10 0 Total volume of cells prepared It would be much appreciated if you can help me wit e recipe of the bumer above. Thank you.
Answer #1

Total cells in the stock suspension=30ml*(7.2*10^6 cells/ml)=216*10^6 cells

To prepare 48 wells number of cells required=1*10^5 cells/well* 48=48*10^5 cells

volume needed to get 48*10^5 cells =48*10^5 cells/7.2*10^6 cells/ml=6.7*10^-1ml=0.7 ml

So 0.7ml of stock is to be taken for dilution

Now V1=volume of stock cell suspension=0.7ml

C1=7.2*10^6 cells/ml

C2=concentration of wells=10^6 cells/ml [1*10^5 cells/100uL=1*10^5 cells/100*10^-6 L*10^3ml/L=1*10^5 cells/0.1ml=10^6cells/ml]

V2=volume of diluted suspension taken for each well

V2=C1 *V1/C2*48=(7.2*10^6 cells/ml)*0.7ml/48*(10^6 cells/ml)=0.1 ml(volume for each well)

Hence 0.7 ml stock is to be diluted to 0.1*48=4.8ml

total volume of cells prepared=4.8ml

0.7 ml stock cells

4.1 ml RPMI added to stock